### v1

@@ 3 The Unbent Bend Number of Plane 4-Graphs
As a new NP-complete version of the SAT problem, we introduce LevelPlanarMonotone2in4Sat, which is defined as follows. As in 2in4Sat, the input is a Boolean formula in conjective normal form where every clause consists of four literals, and the task is to decide whether an assignment of variables to Boolean values exists such that in every clause exactly two literals are true. Additionally, the variable–clause incidence graph is planar and admits a drawing where the variable vertices are placed on a set of parallel lines, called levels, and the clause vertices are placed between the levels such that no edge crosses a level. The levels divide the plane into rectangular strips. The clauses that are represented in the first, third, etc. strip contain only negated literals, whereas the clauses represented in the second, fourth, etc. strip contain only unnegated literals.
@@
Proof (sketch)
We consider only the NP-hardness. Given a LevelPlanarMonotone2in4Sat instance $\Phi$ with $n$ variables, $m$ clauses, and $\ell$ levels, we construct a 4-plane graph that follows the level structure of the variable–clause incidence graph (see Fig. 5 for one level) and has an unbent collection of size 2 with at most $k=2nC$ bends, where $C=2m+\ell$, if and only if $\Phi$ is satisfiable. In terms of flow, the faces representing the clauses (clause gadgets) have a demand of 2, a special face per level has a demand of 1, the external face has a demand of $-C$, and all other faces have a demand of 0. Any path that a unit of flow can take from the external face to one of the faces with positive demand, causes at least $n$ bends. If such a path has exactly $n$ bends, it traverses the variable gadgets. Each variable gadget has two edges as exit doors – one represents assigning this variable to true, and one represents assigning this variable to false. Only one of them can be bent in each of the two drawings of the unbent collection. Each clause gadget is connected to the gadgets of the four variables that occur in the clause. Two units of flow in a clause gadget need to come from two of theses variables gadgets in the one drawing, while, in the other drawing, two units of flow come from the other two variables gadgets. Hence, an unbent collection exists if and only if the exit doors taken by the flow paths correspond to a satisfying truth assignment of $\Phi$.
Next, we describe a constructive 3-approximation algorithm for the unbent bend number. We use the property that the star arboricity of a graph is bounded by its maximum degree.
@@
Otherwise, any optimal unbent collection has size at least 2 and, hence, $\operatorname{OPT}\geq 2b$. We use Proposition 5 to 4-color the edges of $G$ in linear time such that each color induces a star forest. We now define an unbent collection of size 4. For each $i\in[4]$, we construct a drawing $\Gamma_{i}$ where all edges of color $i$ are straight. Each time, we start with $R$ and proceed as follows. Let $e$ be an edge of color $i$, and let $b_{e}=\phi(e)$. If $b_{e}=0$, we do nothing. Otherwise we straighten $e$ by Lemma 3 with respect to the endpoint $v$ of $e$ that is not the center of the star that contains $e$. As a result, the contribution of $e$ to the cost of $\phi$ reduces from $b_{e}$ to 0 and, for each edge $e^{\prime}\neq e$ that is incident to $v$, its contribution to the cost of $\phi$ increases by at most $b_{e}$. Since $v$ has degree at most 4, the cost of the operation is bounded by $-e_{b}+3b_{e}=2b_{e}$. Performing all necessary elementary operations results in a feasible flow $\phi^{\prime}_{i}$ whose cost is at most $\operatorname{cost}(\phi)+\sum_{e\text{ of color }i}2b_{e}$. As the edges of color $i$ form a star forest and we never apply the elementary operation to the center of a star, the operation regarding one edge of color $i$ does not increase the cost of any other edge of color $i$. Thus, using $\phi^{\prime}_{i}$, we can compute an orthogonal representation $R_{i}$ and an orthogonal drawing $\Gamma_{i}$ where all edges of color $i$ are straight. In total, the unbent collection $\{\Gamma_{1},\dots,\Gamma_{4}\}$ has at most $4\operatorname{cost}(\phi)+\sum_{e\in E(G)}2b_{e}=4b+2b=6b$ bends, and the approximation factor of our algorithm is $\operatorname{ALG}/\operatorname{OPT}\leq 6b/(2b)=3$. The algorithm for computing the minimum-cost flow [4] dominates the running time of our algorithm, which hence runs in $O(n^{1+o(1)})$ time.
@@ 4 The Unbent Bend Number of Plane Triconnected Cubic Graphs
In this section we describe how to compute, for a given plane triconnected graph, an unbent collection with the minimum number of bends. Additionally, the collection that we compute consists of only two drawings, which is the minimum.
Let $G$ be a cubic triconnected plane graph. Observe that computing an orthogonal representation of a graph with the minimum number of bends is equivalent to subdividing the edges of the graph the minimum number of times so that the graph admits an orthogonal representation without bends. Hence, our strategy will be to subdivide the edges of $G$ the minimum number of times so that the subdivided graph admits an orthogonal representation with no bends. We call a subdivision vertex a dummy. Let $f_{\operatorname{ext}}$ denote the external face of $G$. Given a cycle $C$, a leg of $C$ is an edge external to $C$ and incident to a vertex of $C$. A $k$-legged cycle $C$ of $G$ is a cycle such that: (i) $C$ has exactly $k$ legs, and (ii) no two legs share a vertex. A $k$-legged cycle with $k\leq 3$ is bad if less than $4-k$ dummies are placed along its edges. Otherwise, it is good. Observe that a 3-legged cycle (such as $C_{5}$ in Fig. 6) is good if we place a dummy along one of its edges, whereas the external cycle is 00-legged and hence it is good if we place four dummies along its edges (which we did in Fig. 6). We use the following lemma.
Lemma 4 ()
[13] A cubic planar graph has an orthogonal representation without bends if and only if it does not contain any bad cycle. Given a cubic planar graph $G$ without bad cycles, an orthogonal representation of $G$ without bends can be computed in linear time.
In Figs. 6(a) and 6(b) we show two ways to add dummies to a cubic planar graph $G$ such that all cycles are good (and such that no edge of $G$ receives dummies in both cases). In the resulting graphs $G^{1}$ and $G^{2}$, every 3-legged cycle contains a dummy and the external cycles contain four dummies. Due to this, both graphs admit orthogonal drawings without bends; see that of $G^{1}$ in Fig. 6(c).
@@
Figure 6: (a–b) Unbent collection $\{G^{1},G^{2}\}$ of a triconnected cubic plane graph $G$ with 3-legged cycles $C_{1},\dots,C_{9}$ and dummies (blue circles). (c) Drawing of $G^{1}$ without bends.
@@
We say that the paths connecting the vertices of a $k$-legged cycle incident to its legs are the contour paths of the cycle. A leg face of a $k$-legged cycle is a face incident to two of its legs. We call a bad cycle of $G$ demanding if it agrees with a coloring rule over a structure that describes the hierarchy of the 3-legged cycles of $G$. This structure and the coloring rule are described in Section 0.C.1. We denote by $D(G)$ denote the set of demanding cycles of $G$ that are pairwise vertex-disjoint. Intuitively speaking, $D(G)$ is a minimum-size set of 3-legged cycles such that, if we place a dummy on an edge of each of them, all 3-legged cycles in $G$ will be good. Some of these dummies can also be shared by the external cycle. For example, in Fig. 6(a), $D(G)=\{C_{3},C_{5},C_{8},C_{9}\}$. We now aim to lowerbound $\operatorname{tbn}(G)$ in terms of the numbers of certain types of demanding cycles.
@@
A cycle $C$ contains another cycle $C^{\prime}$ if the edges of $C^{\prime}$ are edges of the subgraph of $G$ induced by $C$ and its interior vertices. Let $C$ be a demanding cycle. A contour path $P$ of $C$ is interesting if every non-demanding cycle of $G$ that contains $C$ and shares some edges with $C$, contains $P$. For example, $C_{8}$ in Fig. 6(a) has two interesting contour paths, namely those sharing edges with $C_{6}$ and $C_{7}$. See also $C_{5}$ and $C_{9}$, which have only one interesting contour path each, sharing edges with $C_{4}$ and $C_{2}$, respectively.
@@
A demanding cycle $C$ is expensive if and only if $C$ has exactly one interesting contour path $P$ that contains a single edge, and there exists at least a non-demanding 3-legged cycle $C^{\prime}$ such that $C^{\prime}$ contains $C$, shares edges with $P$ and with no other demanding cycle that it contains. For example, in Fig. 6(a), the demanding cycles $C_{5}$ and $C_{9}$ are expensive, as they have one interesting contour path (as observed before) and for both cycles it is composed by one edge. Also, neither $C_{4}$ nor $C_{2}$ shares edges with other demanding cycles that they contain (in particular, $C_{4}$ and $C_{2}$ do not contain other demanding cycles).
@@
A demanding cycle $C$ is short if and only if it has exactly one edge $e$ incident to $f_{\operatorname{ext}}$ and two or more interesting contour paths, one of which is the contour path formed by the single edge $e$. For example, in Fig. 6(a), cycle $C_{8}$ is short, as it has two interesting contour paths (shared with $C_{6}$ and $C_{7}$), one of them is incident to $f_{\operatorname{ext}}$ and composed by one edge.
@@
In Theorem 4.1 we prove that the lower bound of Lemma 5 is tight. Figure 6 shows an unbent collection $\mathcal{C}=\{G^{1},G^{2}\}$ of a cubic planar graph $G$ with $D(G)=\{C_{3},C_{5},C_{8}.C_{9}\}$, $D_{\exp}(G)=\{C_{5},C_{9}\}$, $D_{\operatorname{ext}}(G)=\{C_{3},C_{8},C_{9}\}$, and $D_{\operatorname{short}}(G)=\{C_{8}\}$. Hence, $\mathcal{C}$ has $\operatorname{tbn}(G)=q(G)=2\cdot 4+2+8-(2\cdot 3-1)=13$ dummies.
@@
Proof (sketch)
We sketch how to compute a collection with at most $q(G)$ bends. By Lemma 5, this implies the theorem. We first create two copies of $G$, $G^{1}$ and $G^{2}$, and then we place on them $q(G)$ dummies in total. We prove that, then, none of the two graphs has a bad cycle. By Lemmas 4 and 5, this yields an unbent collection with $\operatorname{tbn}(G)$ bends.
We compute the sets $D(G)$, $D_{\exp}(G)$, $D_{\operatorname{ext}}(G)$, and $D_{\operatorname{short}}(G)$ and we first consider the demanding cycles in $D_{\exp}(G)$. Let $C\in D_{\exp}(G)$. In $G^{1}$, we place a dummy on the only edge belonging to an interesting contour path. In $G^{2}$, we place a dummy on another edge of $C$ (not belonging to an interesting contour path), so that $C$ is good in both $G^{1}$ and $G^{2}$. Still in $G^{2}$, we place a dummy on one of the two legs of $C$ that is contained in all the non-demanding cycles that contain $C$ and share edges with $C$ (such a leg always exists due to Lemma 16 in Section 0.C.3). Hence, for every cycle in $D_{\exp}(G)$, we have placed three dummies; one in $G^{1}$ and two in $G^{2}$. As a result, all demanding cycles in $D_{\exp}(G)$ and all non-demanding cycle sharing edges with them are good in $G^{1}$ and $G^{2}$.
We now consider the cycles in $D(G)\setminus D_{\exp}(G)$. For each such cycle having two or more edges on an interesting contour path, we place a dummy on one such edge in $G^{1}$ and one in $G^{2}$. It remains to place dummies on (i) demanding (non-expensive) cycles that have one edge belonging to an interesting contour path and (ii) on the non-demanding cycles that share edges only with the demanding cycles in case (i). (In Section 0.C.3, we call the non-demanding cycles of case (ii) critical.) To this end, we use a data structure that we call enriched genealogical tree. With its help, we place the dummies that we need to place on such demanding cycles exclusively on interesting contour paths. As a result, all 3-legged cycles become good.

### v2

@@ 3 The Unbent Bend Number of Plane 4-Graphs
As a new NP-complete version of the SAT problem, we introduce LevelPlanarMonotone2in4Sat, which is defined as follows. As in 2in4Sat, the input is a Boolean formula in conjective normal form where every clause consists of four literals, and the task is to decide whether an assignment of variables to Boolean values exists such that, in every clause, exactly two literals are true. Additionally, the variable–clause incidence graph is planar and admits a drawing where the variable vertices are placed on a set of parallel lines, called levels, and the clause vertices are placed between the levels such that no edge crosses a level. The levels divide the plane into rectangular strips. The clauses that are represented in the first, third, etc. strip contain only negated literals, while the clauses represented in the second, fourth, etc. strip contain only unnegated literals.
@@
Proof 3.3 (Proof sketch).
We sketch only the NP-hardness. Given a LevelPlanarMonotone2in4Sat instance $\Phi$ with $n$ variables, $m$ clauses, and $\ell$ levels, we construct a 4-plane graph $G$ that follows the level structure of the variable–clause incidence graph (see Figure 5 for one level) and has an unbent collection of size 2 with at most $k=2nC$ bends, where $C=2m+\ell$, if and only if $\Phi$ is satisfiable.
In $G$, there is a face for each variable vertex (variable gadget; blue in the middle row of Figure 5), for each clause vertex (clause gadget; green and red in the top and bottom row of Figure 5), and for each set of edges connecting a variable vertex to its clause vertices above and, separately, below (tunnel gadget; orange and thin in Figure 5). Each such gadget is surrounded by a set of faces called walls, which make it costly for flow to enter the gadget unless it is through some door. (All walls have “thickness” $2n$, that is, each unit of flow traversing a wall bends at least $2n$ edges.) We construct each face with a specific demand. The clause gadgets have a demand of 2, the level gadgets, which are special faces existing once per level (light purple on the very right in Figure 5), have a demand of 1, the external face has a supply of $C$ (shown as a demand of $-C$ in Figure 5), and all other faces have a demand (and supply) of 0. The demands sum up to $C$.
The blue arrows in Figure 5 indicate valve gadgets. A valve gadget allows a specific amount of flow to be passed from one face (source) to an incident face (target) without creating bends; passing in the reverse direction is only possible with extra bends. Without traversing walls, flow can enter a variable gadget at a door being a valve gadget (on the left side of each variable gadget in Figure 5), and flow can leave a variable gadget at two specific edges (exit doors; red on the top and the bottom side of each variable gadget in Figure 5). Bending the one exit door represents assigning the variable to true, and bending the other exit door represents assigning the variable to false. Each exit door can be bent in only one drawing of the desired unbent collection (of size 2).
Any path that a unit of flow can take from the external face to a face with positive demand, i.e., a clause gadget or a level gadget, causes at least $n$ bends. If such a path has exactly $n$ bends, it traverses variable gadgets and tunnel gadgets, but no walls. Since every variable gadget has one door to enter and two exit doors to leave the gadget (the latter being single edges), all such paths in the first drawing take the same exit door, while all such paths in the second drawing take the other exit door. Observe that the level gadget assures that in both drawings, at least one unit of flow must take such a path starting in the external face and traversing all variable gadgets. The choices of exit doors at the variable gadgets correspond to an assignment of Boolean values to variables. Note that, in Figure 5, there are additional red edges inside the tunnel gadgets. Their purpose is to make any path that a unit of flow can take from the external face to a clause gadget cost $n$ (and not less). Moreover, they assure that the clause gadgets can only be reached via such a tunnel in one of the drawings.
Each clause gadget has four doors, each being a valve gadget allowing one unit of flow to enter. These doors are connected to tunnel gadgets, which, in turn, are connected to the variable gadgets of the four variables that occur in the clause. Two units of flow in a clause gadget need to come from two of these variables gadgets in the one drawing, while, in the other drawing, two units of flow come from the other two variables gadgets. Hence, an unbent collection exists if and only if the exit doors taken by the flow paths correspond to a satisfying truth assignment of $\Phi$.
Next, we describe a 3-approximation algorithm for the unbent bend number. We use the property that the star arboricity of a graph is bounded by its maximum degree. Algor and Alon [1] showed that every $\Delta$-regular graph has star arboricity at most $\Delta/2+O(\Delta^{2/3}(\log\Delta)^{1/3})$. We are interested only in the case $\Delta=4$, and give the following simple explicit construction (stated for general $\Delta$, although it is asymptotically worse than the bound of Algor and Alon).
@@
Otherwise, any optimal unbent collection has size at least 2 and, hence, $\operatorname{OPT}\geq 2b$. We use Proposition 3.4 to 4-color the edges of $G$ in linear time such that each color induces a star forest. We now define an unbent collection of size 4. For each $i\in[4]$, we construct a drawing $\Gamma_{i}$ where all edges of color $i$ are straight. Each time, we start with $R$ and proceed as follows. Let $e$ be an edge of color $i$, and let $b_{e}=\phi(e)$. If $b_{e}=0$, we do nothing. Otherwise we straighten $e$ by Lemma 3.5 with respect to the endpoint $v$ of $e$ that is not the center of the star that contains $e$. As a result, the contribution of $e$ to the cost of $\phi$ reduces from $b_{e}$ to 0 and, for each edge $e^{\prime}\neq e$ that is incident to $v$, its contribution to the cost of $\phi$ increases by at most $b_{e}$. Since $v$ has degree at most 4, the cost of the operation is bounded by $-e_{b}+3b_{e}=2b_{e}$. Performing all necessary elementary operations results in a feasible flow $\phi^{\prime}_{i}$ whose cost is at most $\operatorname{cost}(\phi)+\sum_{e\text{ of color }i}2b_{e}$. As the edges of color $i$ form a star forest and we never apply the elementary operation to the center of a star, the operation regarding one edge of color $i$ does not increase the cost of any other edge of color $i$. Thus, using $\phi^{\prime}_{i}$, we can compute an orthogonal representation $R_{i}$ and an orthogonal drawing $\Gamma_{i}$ where all edges of color $i$ are straight. In total, the unbent collection $\{\Gamma_{1},\dots,\Gamma_{4}\}$ has at most $4\operatorname{cost}(\phi)+\sum_{e\in E(G)}2b_{e}=4b+2b=6b$ bends, and the approximation factor of our algorithm is . The algorithm for computing the minimum-cost flow [16] dominates the running time of our algorithm, which hence runs in $O(n^{1+o(1)})$ time.
@@ 4 The Unbent Bend Number of Plane Triconnected Cubic Graphs
In this section we describe how to compute, for a given plane triconnected cubic graph, an unbent collection with the minimum number of bends. The collections that we compute consist of only two drawings, which is the minimum.
Let $G$ be a plane triconnected cubic graph. Observe that computing an orthogonal representation of a graph with the minimum number of bends is equivalent to subdividing the edges of the graph the minimum number of times so that the graph admits an orthogonal representation without bends. Hence, our strategy will be to compute two different subdivisions of $G$ such that each of the resulting graphs admits an orthogonal representation without bends and the total number of subdivision vertices is minimum. We call a subdivision vertex a dummy. Let $f_{\operatorname{ext}}$ denote the external face of $G$. Given a cycle $C$, a leg of $C$ is an edge external to $C$ and incident to a vertex of $C$. A $k$-legged cycle $C$ of $G$ is a cycle such that: (i) $C$ has exactly $k$ legs, and (ii) no two legs share a vertex. A $k$-legged cycle with $k\leq 3$ is bad if less than $4-k$ dummies are placed along its edges. Otherwise, it is good. Observe that a 3-legged cycle (such as $C_{5}$ in Figure 6) is good if we place a dummy along one of its edges, whereas the external cycle is 00-legged and hence it is good if we place four dummies along its edges (which we did in Figure 6). We use the following lemma.
@@
Figure 6: (a–b) Unbent collection $\{G^{1},G^{2}\}$ of a plane triconnected cubic graph $G$ with 3-legged cycles $C_{1},\dots,C_{9}$ and dummies (blue circles). (c) Drawing of $G^{1}$ without bends.
Lemma 4.1.
[13] A plane cubic graph has an orthogonal representation without bends if and only if it does not contain any bad cycle. Given a plane cubic graph $G$ without bad cycles, an orthogonal representation of $G$ without bends can be computed in linear time.
In Figures 6(a) and 6(b) we show two ways to add dummies to a plane triconnected cubic graph $G$ such that all cycles are good (and such that no edge of $G$ receives dummies in both cases). In the resulting graphs $G^{1}$ and $G^{2}$, every 3-legged cycle contains a dummy and the external cycles contain four dummies. Due to this, both graphs admit orthogonal drawings without bends; see that of $G^{1}$ in Figure 6(c).
@@
We say that the $k$ paths of a $k$-legged cycle connecting the vertices incident to its legs are the contour paths of the cycle. A leg face of a $k$-legged cycle is a face incident to two of its legs. We now aim to lowerbound $\operatorname{tbn}(G)$. To this end, we use specific 3-legged cycles of $G$ that we call demanding. Informally, the demanding cycles form the smallest subset of 3-legged cycles with the property that, if we place a dummy on an edge of each of them, all 3-legged cycles in $G$ will be good. A non-demanding cycle is a 3-legged cycle that is not demanding. For a formal definition of demanding cycles, see the the appendix. The set of demanding cycles is unique. Let $D(G)$ be the largest set of demanding cycles of $G$ that are pairwise vertex-disjoint.
For an example, see Figure 6(a), where $D(G)=\{C_{3},C_{5},C_{8},C_{9}\}$. Observe that all these 3-legged cycles are good due to the placement of four dummy vertices, one for each cycle in $D(G)$. E.g., the dummy vertex on $C_{2}$ also lies on $C_{9}$ (since $C_{2}$ and $C_{9}$ share an edge). Hence, both these 3-legged cycles are good.
The demanding cycles and the set $D(G)$ are computed by using a coloring rule over a structure that describes the hierarchy of the 3-legged cycles of $G$. This structure and the coloring rule are described in the the appendix. We now lowerbound $\operatorname{tbn}(G)$ in terms of the numbers of certain types of 3-demanding cycles.
@@
A cycle $C$ contains another cycle $C^{\prime}$ if the edges of $C^{\prime}$ are edges of the subgraph of $G$ induced by $C$ and its interior vertices. Let $C$ be a demanding cycle. A contour path $P$ of $C$ is interesting if every non-demanding cycle of $G$ that contains $C$ and shares some edges with $C$, contains $P$. For example, $C_{8}$ in Figure 6(a) has two interesting contour paths, namely those sharing edges with $C_{6}$ and $C_{7}$. See also $C_{5}$ and $C_{9}$, which have only one interesting contour path each, sharing edges with $C_{4}$ and $C_{2}$, respectively.
@@
A demanding cycle $C$ is expensive if and only if $C$ has exactly one interesting contour path $P$ that contains a single edge, and there exists at least one non-demanding 3-legged cycle $C^{\prime}$ such that (i) $C^{\prime}$ contains $C$, (ii) $C^{\prime}$ shares edges with $P$, and (iii) $C^{\prime}$ does not share any edges with other demanding cycles that it contains. For example, in Figure 6(a), the demanding cycles $C_{5}$ and $C_{9}$ are expensive, as both have an interesting contour path composed of one edge and neither of the containing cycles ($C_{4}$ and $C_{2}$) contains other demanding cycles.
@@
A demanding cycle $C$ is short if and only if it has exactly one edge $e$ incident to $f_{\operatorname{ext}}$ and two or more interesting contour paths, one of which is the contour path formed by the edge $e$. For example, in Figure 6(a), cycle $C_{8}$ is short, as it has two interesting contour paths (shared with $C_{6}$ and $C_{7}$), one of them is incident to $f_{\operatorname{ext}}$ and composed by one edge.
@@
In Theorem 4.3 we prove that the lower bound of Lemma 4.2 is tight. Figure 6 shows an unbent collection $\mathcal{C}=\{G^{1},G^{2}\}$ of plane triconnected cubic graph $G$ with $D(G)=\{C_{3},C_{5},C_{8}.C_{9}\}$, $D_{\exp}(G)=\{C_{5},C_{9}\}$, $D_{\operatorname{ext}}(G)=\{C_{3},C_{8},C_{9}\}$, and $D_{\operatorname{short}}(G)=\{C_{8}\}$. Hence, $\mathcal{C}$ has $\operatorname{tbn}(G)=q(G)=2\cdot 4+2+8-(2\cdot 3-1)=13$ dummies.
@@
Proof 4.4 (Proof sketch).
We first create two copies of $G$, which we call $G^{1}$ and $G^{2}$. On these, we place $q(G)$ dummies in total. Below we prove that, then, none of the two resulting graphs has a bad cycle. Hence, by Lemma 4.1, they can be drawn without bends, in linear time. Replacing the dummies by bends yields an unbent collection with $q(G)$ bends. Finally, Lemma 4.2 implies the theorem.
We compute the sets $D(G)$, $D_{\exp}(G)$, $D_{\operatorname{ext}}(G)$, and $D_{\operatorname{short}}(G)$. We first consider a demanding cycle $C$ in $D_{\exp}(G)$. In $G^{1}$, we place a dummy on the only edge belonging to an interesting contour path. In $G^{2}$, we place a dummy on another edge of $C$ (not belonging to an interesting contour path), so that $C$ is good in both $G^{1}$ and $G^{2}$. Still in $G^{2}$, we place a dummy on one of the two legs of $C$ that is contained in all the non-demanding cycles that contain $C$ and share edges with $C$ (in the the appendix we prove that such a leg always exists). Hence, for every cycle in $D_{\exp}(G)$, we have placed three dummies; one in $G^{1}$ and two in $G^{2}$. As a result, all demanding cycles in $D_{\exp}(G)$ and all non-demanding cycles sharing edges with them are good in $G^{1}$ and $G^{2}$.
We now consider the cycles in $D(G)\setminus D_{\exp}(G)$. For each such cycle having two or more edges on an interesting contour path, we place a dummy on one such edge in $G^{1}$ and one in $G^{2}$. It remains to place dummies on (i) demanding (non-expensive) cycles that have one edge belonging to an interesting contour path and (ii) on the non-demanding cycles that share edges only with the demanding cycles in case (i). To this end, we use a data structure that we call enriched genealogical tree. With its help, we place the dummies that we need to place on such demanding cycles exclusively on interesting contour paths. As a result, all 3-legged cycles become good.